OpenStudy (anonymous):

An ellipse has the equation (x+3)^2/2^2+(y-7)^2/8^2=1 what are the coordinates of the center of this ellipse? URGENT!!!

ganeshie8 (ganeshie8):

standard equation of ellipse : \[\dfrac{(x-\color{red}{h})^2}{b^2} + \dfrac{(y-\color{red}{k})^2}{a^2} = 1\] center = \((\color{red}{h,~k})\)